# Out of Sight, Still in Mind: Reasoning and Planning about Unobserved Objects with Video Tracking Enabled Memory Models

### v1

@@ IV Experiments & Results
Figure 4: Execution success rate model comparison as a function of (left) the number of objects in the scene, and (right) the number of distractor actions. We find that our approaches perform consistently well across both conditions, outperforming the baseline by a large margin. This is especially prominent in terms of the numbers of distractor actions. The legend applies to both plots.
@@
Objects
4
5
6
7
all
Distractors
1
2
3
all
DOOM
5/5
5/5
5/5
4/5
19/20
DOOM
5/5
4/5
5/5
14/15
LOOM
5/5
4/5
5/5
5/5
19/20
LOOM
4/5
5/5
4/5
13/15
[17, 18]
0/20
[17, 18]
0/15
@@ IV-A Dataset Generation and Relations Definition
We generate a dataset with more than 20,000 skill executions. We generate skill executions with three steps: (1): We create scenes with a variable number of objects in the Isaac Gym simulator [60]. (2): We generate a random skill primitive with random continuous parameters. (3): We execute this skill in simulation. We repeat this process to create a training sequence with multiple timesteps. After the skill executions, we record the point clouds, poses, bounding boxes, and relations pre and post-manipulation. During data collection, if an object $O^{k}_{t}$ disappears at a timestep $t$ due to occlusion, we augment the point clouds at $t$ with the transformed point cloud of the disappeared object from a previous timestep. We show an example of the dataset generation process in Fig. 3. In this section, the number of objects is defined as the number of segments including objects and environments.
@@ IV-B Baseline approach
We use a transformer to implicitly encode the history similar to [19] as our baseline approach. The transformer learns how to combine the history of previous observations and the current observation. For each observation $Z_{t}$, we concatenate the accumulated history with a type embedding $h_{type}$ and feed it into an MLP to get $h_{t}=MLP([E_{p}(E(X_{t}))+h_{t-1},h_{type}(t)])$ while we use $h_{0}=E_{p}(E(X_{0}))$. The type embedding $h_{type}$ encodes whether the segment is observable at the current step. We pass $h_{t}$ to the transformer to model the interaction between object tokens, and then use the decoders to predict object relations and poses.
@@ IV-C Relational prediction evaluation
To test our hypothesis that an explicit object-oriented memory representation is more effective, we evaluate the ability of DOOM and LOOM to predict relations after an action. We define a distractor action as an action applied after some objects disappear from the scene due to occlusion. After an object disappears from the scene, we manipulate other objects several times and then ask the model to reason about the occluded object. In the evaluation across different numbers of objects shown in Table. I, DOOM/LOOM improve over Baseline by close to $20\%$. In the evaluation across different numbers of distractor actions shown in Table. I, DOOM/LOOM improve over Baseline by more than $10\%$. We see that our explicit object-oriented memory approaches consistently outperform the baseline.
IV-D Planning to logical goal relations success rate evaluation
@@ IV-D1 Generalization to a different number of objects
We first show how well our approaches generalize to a variable number of objects. Examples in our training dataset contain a maximum of 5 objects. In the qualitative results shown in Fig. 3, we show that our approaches can generalize to 8 novel objects with different shapes and different viewpoints. In the results shown in Fig. 4, we find that our approaches LOOM and DOOM perform well and are robust to changes in the number of objects presented in the scene. In contrast, the baseline approach performs well on 4 objects, but struggles on scenes with more objects.
@@ IV-D2 Generalization to a different number of distractor actions
The advantage of explicit object-oriented memory is prominent when generalizing to different number of distractor actions (Fig. 4). Our approaches perform well for different numbers of distractor actions, but the baseline performance drops significantly with even one distractor action. We show qualitative results in Fig. 5. Note that distractor actions are not generated during training. This verifies our hypothesis that the performance of the implicit memory approach can significantly decrease with distractor actions. Note that the difference in the planning success rate is more significant than the F1 score of the relational prediction. That is because when evaluating the relational predictions, all relations across all objects were considered, including visible objects, so that the Baseline performs reasonably well. However, when evaluating planning success, the goal relations include occluded objects so the Baseline performs poorly, as it does not reason about occluded objects well.
@@ IV-D3 Real-world planning success evaluation
Since the Baseline performs poorly in simulation evaluation, we only show our approaches DOOM and LOOM in the real-world evaluation. We ran 35 trials per approach with different numbers of objects and numbers of distractor actions. In the results shown in Table II, we find our approaches perform well for different numbers of objects and distractor actions. Note that prior works [17, 18] would achieve a 0% success rate on this evaluation because prior works cannot plan to achieve goals including occluded objects. We show the diversity of the tasks our approaches can achieve in Fig. 1.
@@ IV-E Qualitative Analysis
We visualize several executions of our approaches in Fig. 1, see also our accompanying video. The first example shows how our framework can understand occluded objects and multiple objects with the same appearance. The robot first pushes the yellow mustard to below the shelf. Then the robot pushes the grey cup that is above the shelf as a distractor action. After these two actions, our model can predict the pose of the occluded yellow mustard. The robot can pick and place the blue cup which is above the shelf to achieve the goal relation. For the second example, after the robot puts the apple inside the pitcher, the apple disappears. Then we make the robot move the pitcher as a distractor action. Thanks to the dynamic model with the augmented object inputs, after the distractor action our framework can still understand where is the apple and can pick and place the orange inside the pitcher to make it in contact with the apple to meet the goal relations. For the third example, after the robot picks and places the green box inside the pitcher and moves the pitcher, our framework can understand where the green box is. Then the robot chooses to pick the pitcher and dump it to make the green box in contact with the table. For the fourth example, after the robot pushes the Cheez-It box, two novel objects appear. Our model understands that these two objects are new objects and the Cheez-It box has disappeared, so the robot can achieve the goal of pushing all the objects to the boundary of the table.
@@ V Conclusion
We examined how to encode memory to reason about occluded objects, the appearance of novel objects, and object reappearance. We propose two approaches DOOM and LOOM to incorporate memory into transformer-based relational dynamics. Through comparison to a baseline implicit memory model, we verify that explicit object-oriented memory is better at generalizing in terms of number of objects and number of distractor actions. Our work is the first to jointly leverage a multi-object, relational dynamics model and UVOS object discovery and tracking model to reason about occluded objects. It shows the promise of utilizing object discovery models for long-term manipulation tasks.
@@
We identify several areas for future research. One promising idea is to integrate the tracker and planner and train a full system end-to-end. For example, we could leverage the pose prediction from DOOM as a prior for the tracker. We would additionally like to incorporate a mobile base requiring more complicated memory management by coupling with some form of mapping and examine longer-horizon planning tasks.

### v3

@@ IV Experiments & Results
Figure 4: Execution success rate as a function of (left) the number of objects in the scene, and (right) the number of distractor actions. We find that our approaches perform consistently well across both conditions, outperforming the baseline by a large margin. This is especially prominent in terms of the number of distractor actions. The legend applies to both plots.
@@
Objects 4 5 6 7 all Distractors 1 2 3 all
DOOM 5/5 5/5 5/5 4/5 19/20 DOOM 5/5 4/5 5/5 14/15
LOOM 5/5 4/5 5/5 5/5 19/20 LOOM 4/5 5/5 4/5 13/15
[17, 18] - - - - 0/20 [17, 18] - - - 0/15
@@ IV-A Dataset Generation and Relations Definition
We generate a dataset with more than 20,000 skill executions. We generate skill executions with three steps: (1): We create scenes with a variable number of objects in the Isaac Gym simulator [59]. (2): We generate a random skill primitive with random continuous parameters. (3): We execute this skill in simulation. We repeat this process to create a training sequence with multiple timesteps. After the skill executions, we record the point clouds, poses, bounding boxes, and relations pre and post-manipulation. During data collection, if an object $O^{k}_{t}$ disappears at a timestep $t$ due to occlusion, we augment the point clouds at $t$ with the transformed point cloud of the disappeared object from a previous timestep. We show an example of the dataset generation process in Fig. 3. In this section, the number of objects is defined as the number of segments including object and environment segments.
@@ IV-B Baseline Approach
Our framework is built upon previous works [17, 18], which include comparisons to different baselines and intensive ablation studies. We refer the readers to our prior works to see the effectiveness of our planning framework compared to other alternatives, such as the one mainly relying on explicit pose estimation similar to [31]. In this paper, we use a transformer to implicitly encode the history similar to [19] as our baseline approach. The transformer learns how to combine the history of previous observations and the current observation. For each observation $Z_{t}$, we concatenate the accumulated history with a type embedding $h_{type}$ and feed it into an MLP to get $h_{t}=MLP([E_{p}(E(X_{t}))+h_{t-1},h_{type}(t)])$ while we use $h_{0}=E_{p}(E(X_{0}))$. The type embedding $h_{type}$ encodes whether the segment is observable at the current step. We pass $h_{t}$ to the transformer to model the interaction between object tokens, and then use the decoders to predict object relations and poses.
@@ IV-C Relational Prediction Evaluation
To test our hypothesis that an explicit object-oriented memory representation is more effective, we evaluate the ability of DOOM and LOOM to predict relations after an action. We define a distractor action as an action applied after some objects disappear from the scene due to occlusion. After an object disappears from the scene, we manipulate other objects several times and then ask the model to reason about the occluded object. In the evaluation across different numbers of objects shown in Table. I, DOOM/LOOM improve over Baseline by close to $20\%$. In the evaluation across different numbers of distractor actions shown in Table. I, DOOM/LOOM improve over Baseline by more than $10\%$.
IV-D Planning to Goal Relations Success Rate Evaluation
@@ IV-D1 Generalization to a different number of objects
We first show how well our approaches generalize to a variable number of objects. Examples in our training dataset contain a maximum of 5 objects. In the qualitative results shown in Fig. 3, we show that our approaches can generalize to 8 novel objects with different shapes and different viewpoints. In the results shown in Fig. 4, we find that our approaches LOOM and DOOM perform well and are robust to changes in the number of objects presented in the scene. In contrast, the baseline approach performs well on scenes with 4 objects, but struggles on scenes with more objects.
@@ IV-D2 Generalization to a different number of distractor actions
The advantage of explicit object-oriented memory is prominent when generalizing to different number of distractor actions (Fig. 4). Our approaches perform well for different numbers of distractor actions, but the baseline performance drops significantly with even one distractor action. We show qualitative results in Fig. 5. Note that distractor actions are not generated during training. This verifies our hypothesis that the performance of the implicit memory approach can significantly decrease with distractor actions. Note that the difference in the planning success rate is more significant than the F1 score of the relational prediction. That is because when evaluating the relational predictions, all relations across all objects were considered, including visible objects, so that Baseline performs reasonably well. However, when evaluating planning success, the goal relations include occluded objects so Baseline performs poorly, as it does not reason about occluded objects well.
@@ IV-D3 Real-world planning success evaluation
Since the Baseline performs poorly in simulation, we only show our approaches DOOM and LOOM in the real-world evaluation. We ran 35 trials per approach with different numbers of objects and numbers of distractor actions. In the results shown in Table II, we find our approaches perform well for different numbers of objects and distractor actions. Note that prior works [17, 18] would achieve a 0% success rate on this evaluation because of occluded objects.
@@ IV-E Qualitative Analysis
We visualize executions of our approaches in Fig. 1. The first example shows how our framework can understand occluded objects and multiple objects with the same appearance. The robot first pushes the yellow mustard to below the shelf. Then the robot pushes the grey cup that is above the shelf as a distractor action. After these two actions, our model can predict the pose of the occluded yellow mustard. The robot can pick and place the blue cup above the shelf to achieve the goal relation. For the second example, after the robot puts the apple inside the pitcher, the apple is occluded. Then we make the robot move the pitcher as a distractor action. After the distractor action our framework can still remember the location of the apple and can pick and place the orange inside the pitcher to make it in contact with the apple to meet the goal. For the third example, after the robot picks and places the green box inside the pitcher and moves the pitcher, our framework can understand where the green box is. Then the robot chooses to pick the pitcher and dump it to make the green box in contact with the table. For the fourth example, after the robot pushes the Cheez-It box, two novel objects appear. Our model understands that these are new objects and the Cheez-It box has disappeared, so the robot can achieve the goal of pushing all the objects to the boundary of the table.
@@ V Conclusion
We examined how to encode memory to reason about occluded objects, the appearance of novel objects, and object reappearance. We propose two approaches DOOM and LOOM to incorporate memory into transformer-based relational dynamics. Through comparison to a baseline implicit memory model, we verify that an explicit object-oriented memory is better at generalizing in terms of number of objects and number of distractor actions. Our work is the first to jointly leverage a multi-object, relational dynamics model and a UVOS object discovery and tracking model to reason about occluded objects. It shows the promise of utilizing object discovery models for long-term manipulation tasks.
@@
We identify several areas for future research. One promising idea is to integrate the tracker and planner and train a full system end-to-end. For example, we could leverage the pose prediction from DOOM as a prior for the tracker. We would additionally like to incorporate a mobile base requiring more complicated memory management and examine longer-horizon planning tasks.
Acknowledgments
The authors thank Parker Ewen and Griffin Tabor for the useful discussion. This work was partially supported by NSF Awards #2024778 and #1911232, by DARPA under grant N66001-19-2-4035, by a Sloan Research Fellowship, by ODA award ODA23008GR, and by ONR awards N00024-10-D-6318 and N0014-21-1-2052.
Appendix A Appendix
A-A Preliminary Knowledge
We provide the background information here about the transformers, which is the core component of our approaches. Transformers, an attention-based model, is proposed in  [60] for sequential data. The input for a transformer is a sequence $S$ with length k and the output is another sequence $S^{\prime}$ with the same length k. The heart of the transformer is the attention between different nodes in the sequence $S$. The specific attention used in a transformer is called ”Scaled Dot-Product Attention”. Each input feature is projected to a query, key, and value. The specific function to compute attention is shown in equation. 1. The output feature is a weighted sum of values based on the attention computed from the equation. 1. In this paper, we use the encoder of the transformer with input nodes representing different objects to model the interaction between the objects. For more details about the transformers, we refer you to the original paper [60].
$\texttt{Attention}(Q,K,V)=\texttt{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
A-B Details of UVOS model [28]
We provide implementation details about the UVOS model. In accordance with [28], we have employed Solov2 [61] as a segmentor and STM [62] as a tracker. To familiarize the segmentor with our environment, we fine-tuned it using annotations for both environment segments and objects from YCB-Video dataset [55].
Figure 6: Visualization of objects in the training set for UVOS model.
To be more specific, we annotated a total of $175$ images, which included objects and environments. The environment segments encompass elements such as the robot, table, and shelf. The annotated objects are visually presented in Fig. 6, while objects labeled in the YCB-Video dataset can be found in [55]. To prevent dataset imbalance issues, we down-sampled the YCB-Video dataset before integrating it with the annotated images. As a result, our training dataset comprised $1174$ images and $7105$ object masks. Subsequently, the segmentor was fine-tuned with a learning rate of $61e-$ 6 over $100$ epochs. Note that the segmentor generalizes to novel objects not used during training, as shown in our real-world demonstrations.
In our tasks, key frames in the UVOS approach are designated when the robot is in its home configuration. For each key frame, the UVOS approach employs the segmentor to detect objects within the scene and utilizes the tracker to maintain consistent IDs for previously tracked objects, while those that do not match any previous object are assigned new IDs. During non-key frames, the tracker monitors all segments, relying on its object memory for tracking.
A-C Skill primitives
We define three skill primitives (push, pick-and-place, pick-and-dump) in this paper. We use the same push and pick-and-place skill as [17]. For the details of pick-and-dump skills, we use the same grasp as the pick-and-place skill. After the robot grasps the object, it moves the grasped object to the dump pose. Then the robot rotates the joint 7 by 180 degrees to dump the contained objects. After the dump action, the joint 7 of the robot returns to the original value before the dump action.
Figure 7: Relational prediction F1 score of the different models as a function of the number of objects in the scene.
Figure 8: Relational prediction F1 score of the different models as a function of the number of distractor actions in the scene.
A-D Model details
Figure 9: Two failure cases of our approaches. The first failure case is caused by a failed grasp while the second failure case is caused by unstable placement.
The transformers we used in the paper consist of 2 attention layers, each layer is a 2-head self-attention block. The width of each attention block is 256. The transformer encode takes object features coming from UVOS as input and outputs the object-oriented memory. The dynamics model takes both the object-orientated memory and action tokens as input. For the output of the dynamics model, we discard the action token and only keep the object tokens. For the decoders, the pose decoder $D_{p}$ is a 2-layer MLP with ReLU as an activation function. The relation decoder $D_{r}$ and environment identity decoder $D_{e}$ are 2-layer MLPs with Sigmoid as an activation function.
The input for our model is partial-view point clouds. After the segmentation from UVOS, we use farthest point sampling to downsample each segment to 128 points. The PointConv model consists of 3 layers and outputs a 128-dimension feature for each segment. The learned positional embedding also outputs a 128-dimension feature per ID.
For the details of our Baseline model, $h_{type}$ is a 128 dimension feature per segment to encode whether this segment is observable at the current step. The $MLP$ contains 2 layers with ReLU as an activation function and outputs a 256-dimension feature. We train our approaches and Baseline using the Adam optimizer with $41e-$ 4 as a learning rate.
A-E Limitations
We acknowledge several limitations of the proposed approach. First, we have not integrated the planner and tracker so our framework cannot achieve real-time planning. Second, while we show diverse tasks with different objects like containers and shelves, we have not included demonstrations of more compelling tasks like opening/closing drawers, which would require a robust impedance controller. Third, as shown in the problem definition, our framework assumes a known history of how the objects become occluded, limiting its ability to address occlusion without history.
A-F Failure case analysis
We show two failure cases of our approaches in Fig. 9. For the first example, the robot first puts the apple inside the pitcher and then moves the pitcher a bit. Then the robot receives a goal of putting the orange in contact with the apple. The robot plans to pick the orange but the grasp fails. For the second example, the robot first pushes the yellow mustard below the shelf and then pushes the coffee can. Then the robot picks and places the white cleaner to achieve the goal relation as left(cleaner, yellow mustard) = 1. However, the placement of the cleaner is unstable and thus the goal relation is not successfully achieved. These two failure cases are mainly caused by low-level skill execution failures. The failure cases motivate future directions to implement better low-level skills and incorporate the low-level skills into our high-level planning framework.
A-G Extra simulation results
We show extra visualizations of the comparison to baseline in Fig. 7 and Fig. 8. During the comparison, we find that our approaches DOOM and LOOM consistently outperform the baseline in terms of the relational prediction F1 score.